OpenStudy (anonymous):

help please! i don't know how to use the pressure volume relation in this problem. A fluid system undergoes a nonflow frictionless process following the pressure volume relation as P = (5/V) + 1.5 where P is in bar and V is in m3. During the process the volume changes from 0.15 m3 to 0.05 m3 and the system rejects 45 kJ of heat. Determine: A] change in internal energy B] change in enthalpy

OpenStudy (anonymous):

Are you there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

first find:\[\int\limits_{.15}^{.05} (\frac{ 5 }{ V } +1.5)dV\]

OpenStudy (anonymous):

that will give you the work done...

OpenStudy (anonymous):

seem reasonable?

OpenStudy (anonymous):

delta U = Q-W

OpenStudy (anonymous):

they give you Q so if you have W you know change in internal energy

OpenStudy (anonymous):

you know the conversions?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

alright, i understood now. thanks a lot!

OpenStudy (anonymous):

sure!

OpenStudy (anonymous):

oooooh. really really really thank you. i thought i'm just going to substitue the v's. lol

OpenStudy (anonymous):

heh